Manufacturers Inc. (MI) currently has a labor force of 10, which can produce 600 units per period. The labor cost is now $5000 per period per employee. The company has a long-standing rule that does not allow it to make use of any overtime. In addition, the product cannot be subcontracted, due to the specialized nature of the machinery that MI uses to produce it. As a result, MI can only increase/decrease production by hiring or laying off employees. The cost is $10000 to hire or lay off an employee. Inventory carrying costs are $200 per unit for any unsold items at the end of the period. The inventory level at the beginning of period 1 is 400 units. The forecast demand is 700 in period 1, 600 in period 2, 450 in period 3, 250 in period 4, 500 in period 5, and 550 in period 6.

a.) Compute the costs of the Chase Demand Strategy.

b.) Compute the costs of the Level Strategy

c.) Compare the costs of the two strategies. Which one is superior?
Answer #1
a) Chase demand strategy
With Chase strategy, hiring and layoff is flexible to meet demand and inventory carried forward is minimum
Initial Labor force 10 Labor cost $5,000
Initial production 600 Hire/ layoff cost $10,000
Units produced per labor 60 Inventory carrying cost $200
Period 1 Period 2 Period 3 Period 4 Period 5 Period 6
Forecast Demand 700 600 450 250 500 550
Regular Production 300 600 480 240 480 600
Projected Inventory on hand: 400 300+400-700=0 600-600=0 480-450=30 240+30-250=20 480+20-500=0 600-550=50
Regular Labor Force 5 10 8 4 8 10
Hire 5 4 2
Layoff 10-5=5 2 4
Cost
Period 1 Period 2 Period 3 Period 4 Period 5 Period 6
Regular Labor Cost 5*5000=$25000 $50,000 $40,000 $20,000 $40,000 $50,000
Inventory cost 0 0 30*200=$6000 20*200=$4000 0 50*200=$10000
Hire cost 0 5*10000=$50000 $0 $0 $40,000 $20,000
Layoff cost 5*10000=$50000 $20,000 $40,000 $0 $0
Total $75,000 $100,000 $66,000 $64,000 $80,000 $80,000
$465,000
b) Level Strategy
With level strategy, hiring and layoff is only allowed at beginning and there is level regular production which is constant for all periods.
Level production = (Total demand for 6 periods- Beg. Inventory)/ 6 Level production =(700+600+450+250+500+550-400)/6 =441, so we produce 480 as multiple of labour output per period
Initial Labor force 10 Labor cost $5,000
Initial production 600 Hire/ layoff cost $10,000
Units produced per labor 60 Inventory carrying cost $200
Period 1 Period 2 Period 3 Period 4 Period 5 Period 6
Forecast Demand 700 600 450 250 500 550
Regular Production 480 480 480 480 480 480
Projected Inventory on hand: 400 480+400-700=180 180+480-600=60 60+480-450=90 90+480-250=320 320+480-500=300 300+480-550=230
Regular Labor Force 8 8 8 8 8 8
Hire
Layoff 10-8=2
Cost
Period 1 Period 2 Period 3 Period 4 Period 5 Period 6
Regular Labor Cost 8*5000=$40000 $40,000 $40,000 $40,000 $40,000 $40,000
Inventory cost 180*200=$36000 60*200=$12000 90*200=$18000 320*200=$64000 300*200=$60000 230*200=$46000
Hire cost
Layoff cost 2*10000=$20000
Total $96,000 $52,000 $58,000 $104,000 $100,000 $86,000
$496,000

c)

Clearly cost of Chase strategy ($465,000) is lower than level strategy ($496,000). Hence, Chase strategy is superior.


Please give thumbs up/ likes if you find this answer helpful. Thank you!